Scores on a test can be approximated with a normal distribution. If the average score was an 30 and the standard deviation was 4, find the probability that a randomly chosen test has a score of a) Below 40          b) between 30 and 35



1
Expert's answer
2022-03-16T15:40:40-0400

We have a normal distribution, "\mu=30, \sigma=4."

Let's convert it to the standard normal distribution, "z=\cfrac{x-\mu}{\sigma}."

a) "z=\cfrac{40-30}{4} =2.5;"

"P(X<40)=P(Z<2.5)="

"=0.9938" ​ (from z-table).


b)

"z_1=\cfrac{30-30}{4}=0, \\ z_2=\cfrac{35-30}{4}=1.25,"


"P(30<X<35)=\\P(0<Z<1.25)=\\P(Z<1.25)-P(Z<0)="

"=0.8944-0.5=0.3944" (from z-table).